tan^2 x+tan^2 y+ tan^2 z =1 then what is the maximum value of 3 tanx 4ta y+ 5 tanz?

Grade:12

2 Answers

SanchiSSonawane
5 years ago
Since t
SanchiSSonawane
5 years ago
solution: tanx= tany=tanz=1/√3
then 3/√3 .4/√3 + 5/√3=12/3 + 5/√3=4+5/√3
therefore the answer is 4+5/√3